OpenStudy (anonymous):

freddy and bob are buddies for ever.They love to collect toys. bob has 34 fewer than freddy. altogether they have 200 toys. how many toys does freddy have?

OpenStudy (alexandervonhumboldt2):

b=f-34 b+f=200

OpenStudy (alexandervonhumboldt2):

200-f=f-34 -2f=-166 f=83

OpenStudy (alexandervonhumboldt2):

bob=49 freddi=83

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@davelibe wait its wrong

OpenStudy (anonymous):

its actually right

OpenStudy (alexandervonhumboldt2):

oh yeah

OpenStudy (anonymous):

@AlexandervonHumboldt2 if u add up 49 and 83 it dosent equal up to 200 toys isnt it freddy has134 toys and bob 66?

OpenStudy (alexandervonhumboldt2):

well recount then

OpenStudy (anonymous):

i only needed freddy's